How would you factor:
5x^2-15x-60

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

5x^2-15x-60
= 5(x^2 - 3x - 12)

Answer by stanbon(75887)

How would you factor:
5x^2-15x-60
---
Common Factor: 5
----
= 5(x^2-3x-12)
---
If you want to go further you have to use the Quadratic Formula:
-----
x = [3+-sqrt(9+4*12)]/2
x = [3+-sqrt(57)]/2
x = (3/2)+(1/2)sqrt(57) or x = (3/2)-(1/2)sqrt(57)
--------
Factors would be:
= 5(x-(3+sqrt(57)/2)(x-(3-sqrt(57)/2)
========================================
cheers,
Stan H.